1. fran swims at a speed of 2.2 km/h in still water. the lazy river flows at a speed of 0.6 km/h. how long will it take fran to swim 0.8 km upstream?
2. find the union.
{5,9,11}[union]with the o with slash in the middle
3. a hospital parking lot charges $2.00 for the first hour or part thereof, and $1.00 for each additional hour or part thereof. A weekly pass cost $43.00 and allows unlimited parking for 7 days. if each visit jonny makes to the hospital lasts one and a half hours, what is the minimum number of visits that would make it worthwhile to buy a pass?
4. bayside insurance offers two health plans. under plan a, giselle would have to pay the first 70% of her medical bills, plus 40% of the rest. under plan b, giselle would pay the first $230, but only 25% of the rest. for what amount of medical bills will plan b save giselle money? assume she has over $230 in bills.

Answer by mananth(16946)

1. fran swims at a speed of 2.2 km/h in still water. the lazy river flows at a speed of 0.6 km/h. how long will it take fran to swim 0.8 km upstream?
speed still water 2.2 kmph
river speed = 0.6kmph
upstream speed = 2.2-0.6 = 1.6 kmph
..
time = distance / speed
time = 0.8/1.2
=3/4 hour
..
2. find the union.
{5,9,11}[union]with the o with slash in the middle
3. a hospital parking lot charges $2.00 for the first hour or part thereof, and $1.00 for each additional hour or part thereof. A weekly pass cost $43.00 and allows unlimited parking for 7 days. if each visit jonny makes to the hospital lasts one and a half hours, what is the minimum number of visits that would make it worthwhile to buy a pass?
jonny - 1 1/2 hours /day
first hour $2 + $1 for next half hour.
Total $3 per day
43/3= 14.4 visits
He should make minimum 15 visits to make it worthwhile.
..
4. bayside insurance offers two health plans. under plan a, giselle would have to pay the first 70% of her medical bills, plus 40% of the rest. under plan b, giselle would pay the first $230, but only 25% of the rest. for what amount of medical bills will plan b save giselle money? assume she has over $230 in bills.
let her medical bill be x
plan a - 0.7x +0.3(x-0.7x)
plan b - 230+0.25x
..
0.7x+0.3x -0.21x= 230+0.25x
0.54x=230
x=425.9
Above $ 426 billing plan b will be save her money